OpenStudy (anonymous):

For the function f(x) = 3x+2, find x such that f(x) = 14.

OpenStudy (anonymous):

14 = 3x + 2 find x = ?

OpenStudy (anonymous):

x=4 because 14=3(4)+2 14=12+2 14=14

OpenStudy (anonymous):

That makes sense thank you

OpenStudy (anonymous):

yw